OpenStudy (anonymous):

What is the ratio for the volumes of two similar spheres, given that the ratio of their radii is 3:4? A. 27:64 B. 16:9 C. 9:16 D. 64:27

OpenStudy (michele_laino):

the volume of the first sphere is: \[{V_1} = \frac{{4\pi }}{3}R_1^3\] whereas the volume of the second sphere is: \[{V_2} = \frac{{4\pi }}{3}R_2^3\] so if we divide side by side those formula each other, we get: \[\frac{{{V_1}}}{{{V_2}}} = \frac{{\frac{{4\pi }}{3}R_1^3}}{{\frac{{4\pi }}{3}R_2^3}} = {\left( {\frac{{{R_1}}}{{{R_2}}}} \right)^3}\] now, we have: \[\frac{{{R_1}}}{{{R_2}}} = \frac{3}{4}\] so, please substitute that ratio into the expression for the ratio V_1/V_2, what do you get?

OpenStudy (michele_laino):

formulas*

OpenStudy (anonymous):

@Michele_Laino I'm sorry if you should find be to be incompetent, seeing as how well this was explained, but I do not understand what it is I should be substituting in.

OpenStudy (anonymous):

*me to be

OpenStudy (michele_laino):

it is simple, here is your substitution: \[\Large \frac{{{V_1}}}{{{V_2}}} = \frac{{\frac{{4\pi }}{3}R_1^3}}{{\frac{{4\pi }}{3}R_2^3}} = {\left( {\frac{{{R_1}}}{{{R_2}}}} \right)^3} = {\left( {\frac{3}{4}} \right)^3} = ...?\]

OpenStudy (anonymous):

27/64

OpenStudy (anonymous):

Pretty sure my calculations are right.

OpenStudy (anonymous):

Yep, double checked. Thank you @Michele_Laino! :)

OpenStudy (dan815):

hello there dan :)

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Hey dan :)

OpenStudy (michele_laino):

correct! @Thatsodan